OpenStudy (anonymous):

20/(6-1).2 a.2 b.5 c.8 d.-8 I think the answer is 8 but i don't know.Help

OpenStudy (anonymous):

i got an equation for u... my d*ck + you = a good time

OpenStudy (anonymous):

is that period next to (6-1) and multiplication sign? if yes than First you have to do (6-1)=5 Then: (5).2=10 then: 20/10=2

OpenStudy (anonymous):

thank you :) Forever_Me And yes that is a Multiplication sign.